OpenStudy (anonymous):

Find the inverse of f(x)=4x-3/2 when x=30

OpenStudy (anonymous):

f of x is teckannly y so think of (fx) is another name for y

OpenStudy (anonymous):

I have trouble with the fraction part. How do I get rid of the fraction?

OpenStudy (anonymous):

let me see the fraction please

OpenStudy (anonymous):

4x-3/2

OpenStudy (anonymous):

3 devided by2

OpenStudy (calculusxy):

From what I know, I would insert 30 to x making it to 4(30)-3/2. Seems hard, but looks easy.

OpenStudy (anonymous):

so i substitute 30 for x after I get the inverse?

OpenStudy (calculusxy):

Then u evaluate frm there.

OpenStudy (calculusxy):

Yes beuase x=30 in thi case.

OpenStudy (anonymous):

when i am finding the inverse I get stuck on how to isolate the y variable, because it ends up in a fraction.

OpenStudy (calculusxy):

Th inverse when factoring js multiplying.

OpenStudy (anonymous):

so i just multiply the fraction to get rid of it?

OpenStudy (jdoe0001):

\(\bf f(x)=y=4x-\cfrac{3}{2}\qquad inverse \implies x= 4y-\cfrac{3}{2}\\ \quad \\ \quad \\ x= 4y-\cfrac{3}{2}\implies x+\cfrac{3}{2}=4y\implies \cfrac{\left(x+\frac{3}{2}\right)}{4}=y\\ \quad \\ \cfrac{x}{4}+\cfrac{\frac{3}{2}}{4} = y\implies \cfrac{x}{4}+\cfrac{3}{2}\cdot \cfrac{1}{4} = y\implies \cfrac{x}{4}+\cfrac{3}{8} = y = f^{-1}(x)\)

OpenStudy (anonymous):

The fraction is set up like this in the problem

OpenStudy (anonymous):

how would i find the inverse of that?

OpenStudy (calculusxy):

Yes you multiply

OpenStudy (jdoe0001):

\(\bf f(x)=y=\cfrac{4x-3}{2}\qquad inverse \implies x= \cfrac{4y-3}{2}\\ \quad \\ \quad \\ 2x = 4y-3\implies 2x+3 = 4y\implies \cfrac{2x+3}{4} = y = f^{-1}(x)\)

OpenStudy (anonymous):

OOh, I see it now XD thnx!

OpenStudy (jdoe0001):

yw